A segment 35cm long is divided into 2 segments whose ratio is 3:4
How long is each segment?

Answer by Cromlix(4381)

Ratio 3:4
Add 3 + 4 = 7
3/7 * 35 = 15cm
4/7 * 35 = 20cm
Hope this helps.
:-)